OpenStudy (anonymous):

. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this. -x + 2y = 8 -x + 10y = 48

OpenStudy (anonymous):

subtract the first equation from the second equation to get: 8y = 40 can you take it from here?

OpenStudy (anonymous):

Do you know how to use Matrices?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Ok use dplanc's approach.

OpenStudy (anonymous):

HOW WOULD I DO

OpenStudy (anonymous):

change the signs of each term in the first equation to get x-2y = -8 Then add the equations x - 2y = -8 -x + 10y = 48 --------------- 0x + 8y = 40 So we have 8y = 40 Solve 8y = 40 for y WHAT WILL I DO NEXT

OpenStudy (callisto):

-x + 2y = 8 -(1) -x + 10y = 48 -(2) (2) -(1) (-x + 10y) - (-x + 2y) = 48 - 8 -x + 10y + x - 2y = 40 8y = 40 y = ...? Sub y = ... into (1) to get x.

OpenStudy (anonymous):

SO WHERE THERE IS A Y PUY 1

OpenStudy (callisto):

Can you solve y first 8y = 40 y = ...?

OpenStudy (anonymous):

Y=5

OpenStudy (anonymous):

If you have two equations with two unkowns (x and y) , one way to solve this "system" of equations is to multiply one equation by a constant such that when it is added to the second one of the variables is eliminated from the resulting equation. for example for the system of equations \[2x+3y=4 \] with \[x+y=3\] we can multiply the second equation by \[-2\] to get \[-2x-2y=-6\], you are then allowed to add this to the first equation; \[ (2x+3y)+(-2x-2y) =4-6 \] gives \[ 0x+(3-2)y=-2\] and \[y=-2\] once you have this substitute it back into equation \[2x+3y=4\] and get \[2x-6=4\] then solveing for x, \[2x=-2\],\[x=-1\].

OpenStudy (anonymous):

Use the same idea for your problem.

OpenStudy (anonymous):

Or do you need another example?

OpenStudy (callisto):

As you've found, y=5 Now, put y=5 into (1) -x + 2(5) = 8 Can you find x?